OpenStudy (anonymous):

3x^2-x-2 show all work please

OpenStudy (anonymous):

Use the AC factoring method (no guessing). Find 2 numbers b1 and b2 that satisfy: Product: b1*b2 = a*c = -6; and sum (b1 + b2) = -1. Replace in the equation the term (-x) by the 2 terms (-x) by the 2 terms (x) and (-2x), then factor.